OpenStudy (anonymous):

Simplify (6y2 -6/ 8y2 +8y)/(3y -3 / 4y2 + 4y) Help? @campbell_st @IMStuck

OpenStudy (campbell_st):

well 1st thing is ... flip the 2nd fraction and multiply \[\frac{6y^2 -6}{8y^2 + 8y} \times \frac{4y^2 + 4y}{3y - 3}\] do some factoring \[\frac{6(y + 1)(y - 1)}{2(4y^2 + 4y)} \times \frac{(4y^2 + 4y)}{3(y - 1)}\] now remove common factors for the answer...